# Assessing Open-world Forgetting in Generative Image Model Customization

### v1

@@ 4.2 Appearance Drift
Figure 3: Appearance drift as consequence of customization. a) chromaticity plot of pixels of three realization of the prompts (’photo of a car/cow’) and the same seed with different models, namely b) the base model, c) model adapted to lighthouse and d) model adapted to bike.
@@
We evaluate the appearance drift using the CDI together with KID (FID results are presented in Appendix B.3). We conducted a comprehensive experiment adopting the carefully curated selection of common concepts from the dataset of Torralba & Oliva (2003). For each prompt, we generated 1,000 images using both original and the ten models from the Customized Model Set. Figure 4 presents the mean values of the metrics across several adaptations, providing a visual representation of the differences captured by each measure. For a detailed overview of the results, including individual model performances, refer to Table 8 in the appendix. To help interpret the KID and CDI scores, we provide a control setting (blue line). In this configuration, we measure CDI and KID between images generated with the base model but using different seeds (functioning as a lower bound). If we are sampling from the same distribution, the base model should yield lower distances (approaching zero as the number of samples grows) than the customized models (DB and CD).
@@ 4.3 Local Drift
Figure 5: Similarity (measured as cosine distance in CLIP-I embedding space) and perceptual metrics between models before and after adaptation. For each concept trained, we evaluate closely related concepts to measure the local drift. a) Results with DreamBooth adaptation (includes prior regularization). b) Results with DreamBooth with Drift Correction. c) Color Drift Index (CDI) and Kernel Inception Distance (KID). For more results see Appendix A.
Here, we repeated our experiments from Section 4.1 and 4.2 to measure the local semantic and local appearance drift. For this setup, we generated 1,000 samples of the closest concepts (superclasses) to each trained model (see Appendix B.1 for the details) and evaluated the CLIP-I, CDI, FID, and KID metrics. As Figure 5a shows, the semantic drift is showing a significant shift towards the left, indicating that local drift is more pronounced. For appearance drift, Figure 5c depicts a more uniform color and KID shift over all the models; this shows that related concepts are affected with a similar magnitude by appearance drift. Again the application of our proposed Drift Correction method greatly reduces both the local semantic drift (as measured in Figure 5b and it almost removes the local appearance drift as measured by KID, even though some color drift remains (see Figure 5c).
@@ 5 Discussion and Conclusion
Our investigation into unintended consequences of generative model adaptation reveals several key contributions: 1) We observe that finetuning foundational generative models leads to substantial open-world forgetting, manifesting as both semantic and appearance drift. 2) We propose to measure semantic drift by considering zero-shot classification on a wide range of image classification tasks, and we propose a novel metric to measure color drift that combined with KID measures appearance drift. 3) We propose a technique to mitigate open-world forgetting using functional regularization. This method shows promise in preserving foundational knowledge while allowing effective customization. These findings provide a guidance and initial step to tackle open-world forgetting and developing more robust and reliable customization techniques.
The last couple of years has seen a growing number of foundation models being released. These models have been enthusiastically adapted by the community, and are often the basis for further finetuning operations to new domains, tasks, etc. The main purpose of our paper is to contribute to the awareness of the unintended changes in the original foundation model knowledge that can occur when adapting these. This phenomenon, coined open-world forgetting, is complicated to assess, and we have proposed several ways to assess it for customized generative models. We hope that our results motivate more research in this direction. Open-world forgetting resembles the finding of a needle in a haystack due to the vast amount of prior knowledge of the foundation models, and we think that other research directions which actively optimize to find the most affected areas might be required. Finally, we consider extending our methodology to other forms of model adaptation, such as unlearning techniques as exiting future work.
Limitations. While our study provides valuable insights into open-world forgetting in generative model adaptation, there are several limitations to consider. First, inherently to the nature of open-world, we cannot exhaustively evaluate the entire conceptual space of the model. Our analysis is limited to a subset of concepts and may not capture all potential instances of forgetting. Second, our study has mainly focused on evaluating the impact of diffusion model customization. Even though we represent a mitigation technique for open-world forgetting (Drift Correction) we think that there are many more ways to address this problem (especially considering the vast continual learning literature) which have not considered in our study.
@@
Appendix A More examples
A.1 Open-world forgetting
A.2 Local drift
In Figure 7 we present examples of local drift for the concept ”dog”.
Appendix B Experiment details
@@
B.1 Semantic Drift Evaluation
@@
To evaluate open-world forgetting, we select a random subset of 10,000 user prompts from DiffusionDB 2M (Wang et al., 2023) (Table 3). For adaptation training and evaluation, we choose a subset of 10 concepts from CustomConcept101 (Kumari et al., 2023), namely decoritems_vase2, decoritems_lamp1, person_2, person_3, pet_cat5, pet_dog4, transport_bike, transport_car2, scene_lighthouse, scene_waterfall. Each concept contains approximately 3-5 images. For superclass evaluation, we create a dataset of 10 synonyms with respect to each concept, which can be found in Table 4.
@@ B.2 Diffusion Classifier
The standard deviations from Table 1 and 2 are included in Table 6 and 7, respectively.
Food CIFAR10 Aircraft Pets Flowers STL10 ObjectNet
Base Model 71.09 81.60 23.40 86.87 50.00 93.00 28.50
[…]
Table 7: Concept fidelity (DINO, CLIP-I) and prompt fidelity (CLIP-T). Drift Correction maintains fidelity across metrics.concept evaluation
DINO CLIP-I CLIP-T
DB 0.4241$\pm$0.1503 0.6764$\pm$0.1046 0.7896$\pm$0.0296
DB-DC 0.4283$\pm$0.1584 0.6817$\pm$0.1097 0.7799$\pm$0.0324
[…]
CD 0.4422$\pm$0.1378 0.6934$\pm$0.0902 0.7916$\pm$0.0266
CD-DC 0.4381$\pm$0.4381 0.6925$\pm$0.0888 0.7899$\pm$0.0280
[…]
B.3 Appearance drift full tables
See Table 8 for the full results which have been used for the generation of Figure 4. The prompt are ‘0:Face’,‘1:Pedestrian’,‘2:Car’,‘3:Cow’,‘4:Hand’,‘5:Chair’,‘6:Mountain’, ‘7:Beach’,‘8:Forest’,‘9:Highway’,‘10:Street’,‘11:Indoor’, ‘12:Animal in natural scene’, ‘13:Tree in urban scene’, ‘14:Close-up person in urban scene’,‘15:Far pedestrian in urban scene’,‘16:Car in urban scene’,‘17:Lamp in indoor scene’, ‘18:empty prompt’.
Prompt Vanilla DB DB-DC CD CD-DC
CDI FID KID CDI FID KID CDI FID KID CDI FID KID CDI FID KID
Appendix C Ablations
C.1 Does finetuning lead to loss of diversity?
Our proposed method not only improves the mitigation of subject shifting, as evidenced in Figures 4 and 5, but also maintains the diversity of the original model. To validate this, we conducted an extensive evaluation using 100 different prompts, each generating 100 images. These prompts were sourced from the DiffusionDB subset, as detailed in Appendix B.1.
C.2 Increasing the buffer size reduces drift
Table 9: Effect of the number of images in the buffer.
Metric 0 50 100 200 500 1000 2000
CDI 0.87 0.64 0.58 0.56 0.70 0.60 0.51
C.3 Influence of the training images

### v2

@@ 4.2 Appearance Drift
Figure 3: Appearance drift as consequence of DreamBooth customization. a) chromaticity plot of pixels of three realization of the prompts (’photo of a car/cow’) and the same seed with different models, namely b) the base model, c) model adapted to lighthouse and d) model adapted to bike.
@@
We evaluate the appearance drift using the CDI together with KID (FID results are presented in Appendix B.2). We conducted a comprehensive experiment adopting the carefully curated selection of common concepts from the dataset of Torralba & Oliva (2003). For each prompt, we generated 1,000 images using both original and the ten models from the Customized Model Set. Figure 4 presents the mean values of the metrics across several adaptations, providing a visual representation of the differences captured by each measure. To help interpret the KID and CDI scores, we provide a control setting (blue line). In this configuration, we measure CDI and KID between images generated with the base model but using different seeds (functioning as a lower bound). If we are sampling from the same distribution, the base model should yield lower distances (approaching zero as the number of samples grows) than the customized models.
@@
For a detailed overview of the results, including individual model performances and additional customization methods, refer to Tables 4 and 5 in Appendix B.2.
@@
User Study
We conducted a user study to verify the effectiveness of the proposed drift correction to mitigate open-world forgetting. We evaluate over 100 image pairs (before/after customization) for both DreamBooth and Custom Diffusion with and without drift correction performed with 24 participants across 20 concepts for both DreamBooth and Custom Diffusion. The results strongly favor our approach, with participants preferring our drift-corrected outputs 66% of the time compared to baseline customization. The statistical analysis confirms the significance of these results (paired t-test: $t(23)=19.624$, $p<0.001$; Wilcoxon signed-rank test: $p<0.001$). We have included the settings and sample images of the user study in Appendix D.
@@ 4.3 Local Drift
Figure 5: Similarity (measured as cosine distance in CLIP-I embedding space) and perceptual metrics between models before and after adaptation. For each concept trained, we evaluate closely related concepts to measure the local drift. a) Results with DreamBooth adaptation (includes prior regularization). b) Results with DreamBooth with Drift Correction. c) Color Drift Index (CDI) and Kernel Inception Distance (KID). For more results see Appendix F.
Here, we repeated our experiments from Section 4.1 and 4.2 to measure the local semantic and local appearance drift. For this setup, we generated 1,000 samples of the closest concepts (superclasses) to each trained model (see Appendix E.1 for the details) and evaluated the CLIP-I, CDI, FID, and KID metrics. As Figure 5a shows, the semantic drift is showing a significant shift towards the left, indicating that local drift is more pronounced. For appearance drift, Figure 5c depicts a more uniform color and KID shift over all the models; this shows that related concepts are affected with a similar magnitude by appearance drift. Again the application of our proposed Drift Correction method greatly reduces both the local semantic drift (as measured in Figure 5b and it almost removes the local appearance drift as measured by KID, even though some color drift remains (see Figure 5c).
@@ 5 Discussion and Conclusion
Our investigation into unintended consequences of generative model adaptation reveals several key findings. First, we demonstrate that finetuning foundational generative models leads to substantial open-world forgetting, manifesting as both semantic and appearance drift. Our results show that even minor adaptations can cause significant deterioration in the model’s ability to maintain its original capabilities. To quantify these effects, we introduced novel evaluation approaches: measuring semantic drift through zero-shot classification performance across diverse image classification tasks, and assessing appearance drift through our proposed Color Drift Index combined with traditional metrics like KID. These methods provide a framework for understanding and measuring the impact of model adaptation on both semantic understanding and visual representation. Additionally, we propose a technique to mitigate open-world forgetting using functional regularization. Our experiments demonstrate that this method effectively preserves foundational knowledge while allowing for successful customization, offering a promising direction for developing more robust adaptation techniques.
We have focussed our analysis on the forgetting of customization methods, however, the adaptation of diffusion models occurs in various domains. The understanding and addressing of open-world forgetting is also important in these fields, and we believe the developed metrics can be useful. As a preliminary analysis, we have included results in Appendix A.1 for unlearning, where concepts are removed from a diffusion model, and quantization, where quantization is applied to obtain a more compute-efficient version of the model. The initial results confirm that indeed these operations lead to forgetting, and our metrics can be used to study this phenomenon. Another future research direction which we aim to explore is the development of optimization methods to identify the most affected areas in the latent space during adaptation.
@@
Appendix A Additional Experiments
A.1 Analysis of Open-World Forgetting Across Model Modifications
To demonstrate that open-world forgetting is a general phenomenon that occurs across different types of model modifications, we analyzed three broad categories of model alterations: model customization, unlearning, and quantization methods.
For customization, we evaluated DreamBooth Ruiz et al. (2023), Custom Diffusion Kumari et al. (2023b), and BOFT Liu et al. (2024), which are commonly used for personalizing generative models, learning decoritems_lamp1 concept. For unlearning, we examined NSFW and nudity content removal approaches including Erasing Concepts Gandikota et al. (2023) and Ablating Concepts Kumari et al. (2023a), as well as RECE Gong et al. (2025). Finally, for quantization, we compared LLM.int8() Dettmers et al. (2022), NF4 from QLORA Dettmers et al. (2024), and SVDQuant Li et al. (2025) methods. All similarities were computed using 10,000 DiffusionDB prompts, similar to Figure 2.
As shown in Figure 6, all modification approaches exhibit significant distributional shifts from the base model’s behavior, showing similar distributions. This suggests comparable levels of forgetting despite their different adaptation approaches, with SVDQuant showing a particularly sharp peak near similarity 0.95, indicating more consistent but still present forgetting effects. These results reinforce our hypothesis that open-world forgetting is inherent to weight modification, regardless of the intended purpose or method of modification.
(a) Customization
(b) Unlearning
(c) Quantization
Figure 6: Distribution of CLIP-I cosine similarities between base model and modified model outputs across different modification approaches. All approaches exhibit significant and similar shifting behavior.
A.2 Large-scale Concept Evaluation
To assess the generalizability of our findings, we conducted an experiment using all concepts from CustomConcept101. We evaluated zero-shot classification performance through multiple random sampling runs for each concept count N (1 to 100), computing confidence intervals for statistical reliability. For each dataset and N value, we performed 10 independent sampling runs.
Figure 7 shows the mean classification score in seven datasets (Aircraft, CIFAR-10, Flowers, Food, ObjectNet, Pets, and STL-10), plotting the mean classification score against the number of concepts on a logarithmic scale. Demonstrating consistent classification performance across different scales, persisting across datasets with varying baseline performance levels, from STL-10 ($\sim 90\%$) and Pets ($\sim 80\%$) to Aircraft ($\sim 20\%$).
The nearly horizontal trend lines and narrow confidence intervals suggest our findings about representation drifting are scale-invariant. This implies that insights from smaller-scale evaluations can reliably generalize to larger concept sets.
Figure 7: Zero-shot classification performance as a function of the number of concepts used in evaluation. The stability across all datasets suggests that semantic drift findings generalize beyond the small-scale evaluations.
A.3 Cross-Architecture Analysis
To investigate whether open-world forgetting is specific to particular architectures, we applied DreamBooth to different families of generative models: Stable Diffusion 1.5 Rombach et al. (2022), Stable Diffusion 3.5 Medium Esser et al. (2024), FLUX.1 [dev] Labs (2023), and Sana 1.6B Xie et al. (2024). Despite their architectural differences, all models exhibit similar distributional shifts when modified through DreamBooth (Figure 8). The rightward skew in the similarity distributions suggests that while the degree of forgetting may vary slightly between architectures, the fundamental phenomenon persists across model families. This provides evidence that open-world forgetting emerges naturally from weight adjustments in large-scale generative models, rather than being a byproduct of particular architectural decisions.
Figure 8: CLIP-I similarity distributions between base and DreamBooth-modified versions across different model architectures. All show characteristic shifts in their output distribution.
Appendix B Additional Results
B.1 Semantic Drift
In this section we show additional methods for Table 1 as well as standard deviations for statistical significance.
Food CIFAR10 Aircraft Pets Flowers STL10 ObjectNet
Base Model 71.09 81.60 23.40 86.87 50.00 93.00 28.50
[…]
SVDiff 77.40$\pm$2.32 91.45$\pm$1.72) 46.94$\pm$1.29 84.66$\pm$1.01 25.88$\pm$1.95 65.82$\pm$2.78 19.32$\pm$0.76
SVDiff-DC 80.60$\pm$1.40 92.95$\pm$0.50 48.35$\pm$1.10 85.69$\pm$0.71 26.77$\pm$1.39 69.27$\pm$1.62 20.04$\pm$0.68
BOFT 76.60$\pm$4.76 89.75$\pm$5.20 46.04$\pm$2.39 82.86$\pm$1.93 25.26$\pm$2.10 64.79$\pm$1.90 20.05$\pm$1.98
BOFT-DC 78.72$\pm$1.29 91.20$\pm$1.75 48.43$\pm$1.90 85.44$\pm$1.17 26.20$\pm$1.54 67.33$\pm$0.58 19.35$\pm$1.84
DINO CLIP-I CLIP-T
DB 0.4241$\pm$0.1503 0.6764$\pm$0.1046 0.7896$\pm$0.0296
DB-DC 0.4283$\pm$0.1584 0.6817$\pm$0.1097 0.7799$\pm$0.0324
[…]
CD 0.4422$\pm$0.1378 0.6934$\pm$0.0902 0.7916$\pm$0.0266
CD-DC 0.4381$\pm$0.4381 0.6925$\pm$0.0888 0.7899$\pm$0.0280
[…]
SVDiff 0.5017$\pm$0.1404 0.7493$\pm$0.0578 0.7442$\pm$0.0497
SVDiff-DC 0.4288$\pm$0.1804 0.6960$\pm$0.1144 0.7287$\pm$0.0470
BOFT 0.3765$\pm$0.1797 0.6674$\pm$0.0983 0.7547$\pm$0.0291
BOFT-DC 0.4043$\pm$0.1789 0.6705$\pm$0.0991 0.7618$\pm$0.0341
B.2 Appearance Drift
See Tables 4 and 5 for the full results which have been used for the generation of Figure 4. The prompt are ‘0:Face’,‘1:Pedestrian’,‘2:Car’,‘3:Cow’,‘4:Hand’,‘5:Chair’,‘6:Mountain’, ‘7:Beach’,‘8:Forest’,‘9:Highway’,‘10:Street’,‘11:Indoor’, ‘12:Animal in natural scene’, ‘13:Tree in urban scene’, ‘14:Close-up person in urban scene’,‘15:Far pedestrian in urban scene’,‘16:Car in urban scene’,‘17:Lamp in indoor scene’, ‘18:empty prompt’.
Table 4: Comparison of CDI, FID, and KID Values for custom vs custom-regularized methods (similar to those in Figure 4).
Method CDI $\downarrow$ KID $\downarrow$ FID $\downarrow$
Base Model 0.11$\pm$0.03 0.01$\pm$0.01 31.32$\pm$10.15
DB 0.56$\pm$0.19 1.41$\pm$0.82 45.97$\pm$12.39
DB-DC 0.30$\pm$0.11 0.27$\pm$0.19 29.07$\pm$7.59
CD 0.52$\pm$0.13 1.40$\pm$0.70 46.91$\pm$12.89
CD-DC 0.30$\pm$0.07 0.42$\pm$0.35 33.06$\pm$12.26
SVDiff 0.87$\pm$0.14 1.15$\pm$0.67 52.55$\pm$12.95
SVDiff-DC 0.30$\pm$0.03 0.39$\pm$0.22 45.64$\pm$12.74
BOFT 0.57$\pm$0.11 1.12$\pm$0.72 52.66$\pm$18.94
BOFT-DC 0.34$\pm$0.05 0.36$\pm$0.20 40.82$\pm$10.04
Prompt Vanilla DB DB-DC CD CD-DC
CDI FID KID CDI FID KID CDI FID KID CDI FID KID CDI FID KID
Appendix C Ablations
C.1 Does finetuning lead to loss of diversity?
Our proposed method not only improves the mitigation of subject shifting, as evidenced in Figures 4 and 5, but also maintains the diversity of the original model. To validate this, we conducted an extensive evaluation using 100 different prompts, each generating 100 images. These prompts were sourced from the DiffusionDB subset, as detailed in Appendix E.1.
C.2 Increasing the buffer size reduces drift
Table 6: Effect of the number of images in the buffer.
Metric 0 50 100 200 500 1000 2000
CDI 0.87 0.64 0.58 0.56 0.70 0.60 0.51
C.3 Influence of the training images
Appendix D User Study
We conducted a user study with 24 participants, evaluating 100 image pairs (before/after customization) across 20 concepts for both DreamBooth and Custom Diffusion. In each trial, participants were asked “Which image is the most similar to the center image?” where the center image was generated by the base model, and the left and right images (randomly assigned) were from the customized model and the customized model with our drift correction method.
In the user study, participants were presented with image triads, as shown in Figure 12. Each triad consisted of a reference image in the center and two comparison images (labeled A and B) on either side. The methods evaluated were DreamBooth and Custom Diffusion and its corresponding Drift Corrected versions. Participants were given the following instructions:
”Look at the three images shown: one in the center, and two options (A and B) on the sides. Your task is to determine which side image (A or B) is more visually similar to the center image.”
The results strongly favor our approach, with participants preferring our drift-corrected outputs 66% of the time compared to baseline customization. The standard deviation across users was just 3.95%, and statistical analysis confirms the significance of these results (paired t-test: $t(23)=19.624$, $p<0.001$; Wilcoxon signed-rank test: $p<0.001$). These results demonstrate consistent preference for our method across all participants and validate our quantitative metrics with human perception.
Figure 12: Example triad presented in the user study.
Appendix E Experiment Details
@@
E.1 Semantic Drift Evaluation
@@
To evaluate open-world forgetting, we select a random subset of 10,000 user prompts from DiffusionDB 2M (Wang et al., 2023) (Table 7). For adaptation training and evaluation, we choose a subset of 10 concepts from CustomConcept101 (Kumari et al., 2023b), namely decoritems_lamp1, decoritems_vase2, person_2, person_3, pet_cat5, pet_dog4, transport_bike, transport_car2, scene_lighthouse, scene_waterfall. Each concept contains approximately 3-5 images. For superclass evaluation, we create a dataset of 10 synonyms with respect to each concept, which can be found in Table 8.
@@ E.2 Diffusion Classifier
Appendix F Qualitative Results
F.1 Open-world forgetting
F.2 Local drift
In Figure 14 we present examples of local drift for the concept ”dog”.
F.3 Comparison with Drift Correction
To demonstrate the effectiveness of our proposed correction method, we provide visual examples showing how the pretrained model experiences semantic and appearance drifting, and how our method mitigates these issues. The comparative results are presented in Figures 15 and 16. These examples clearly illustrate both types of drifting in the baseline model and the improvements achieved through our correction approach.
Figure 15: Qualitative result comparisons on diverse prompts for the pretrained model, a customization method and the proposed Drift Correction. These are random results and all generated from the same initial seed.
Figure 16: More qualitative result comparisons on diverse prompts for the pretrained model, a customization method and the proposed Drift Correction. These are random results and all generated from the same initial seed.